OpenStudy (anonymous):

Evaluate the expression when a = -1.4 and b = -2.7. |a| + b

OpenStudy (anonymous):

-1.3 i guess

OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A.) 5x - 8y = 9 12x - 5y = 7 B.) 3x + 6y = 8 y = 3x - 8 C.) y = 2x - 7 x + 3y = 3 D.) 2y = -3x + 5 2x - 3y = 4

OpenStudy (anonymous):

option D